OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x + 3y = -45.5 -5x + 8y = -43

OpenStudy (anonymous):

substitute/eliminate x = -3/4y-91/8 substitute/eliminate y = -17/2 Solution: x = -5; y = -17/2; (x, y) = (-5, -8.5)